### v2

@@ Abstract
Current deep regression models usually learn in point-wise way that treat each sample as an independent input, neglecting the relative ordering among different data. Consequently, the regression model could neglect the data ’s interrelationships, potentially resulting in suboptimal performance. Moreover, the existence of aleatoric uncertainty in the training data may drive the model to capture non-generalizable patterns, contributing to increased overfitting. To address these issues, we propose a novel adaptive pairwise learning framework (AdaPRL) for regression tasks which leverages the relative differences between data points and integrates with deep probabilistic models to quantify the uncertainty associated with the predictions. Additionally, we adapt AdaPRL for applications in multi-task learning and multivariate time series forecasting. Extensive experiments with several real-world regression datasets including recommendation systems, age prediction, time series forecasting, natural language understanding, finance, and industry datasets show that AdaPRL is compatible with different backbone networks in various tasks and achieves state-of-the-art performance on the vast majority of tasks, highlighting its notable potential including enhancing prediction accuracy and ranking ability, increasing generalization capability, improving robustness to noisy data, improving resilience to reduced data and enhancing interpretability. Experiments also shows that AdaPRL can also be seamlessly incorporated into recent proposed regression frameworks.
@@ 1 Introduction
In recent years, regression tasks have become increasingly important in various domains, including computer vision, time series forecasting, recommendation systems, finance, and so on. Traditional regression methods which integrate with point-wise loss function such as mean squared error (MSE), mean absolute error (MAE) and Huber have been widely adopted due to the simplicity and effectiveness. However, these traditional pointwise methods often struggle with capturing the nuanced relationships between data points, potentially resulting in less effective performance. Pairwise learning, on the other hand, offers a promising alternative by focusing on relative differences between pairs of data points rather than absolute values.
@@
To overcome the limitations of point-wise learning, Learning-to-Rank (LTR) approaches have been introduced, where pairwise methods aim to ensure that the predicted score of a positive example surpasses that of a negative sample for each positive-negative pair, and listwise methods promote higher rankings for positive samples across the entire list of items. Representative pairwise learning methods include RankSVMJoachims (2002), GBRankZheng et al. (2007), RankNetBurges et al. (2005), LambdaRankBurges et al. (2006) and LambdaMARTBurges (2010). Some previous studies Li et al. (2015); Yan et al. (2022); Yue et al. (2022); Sheng et al. (2023) have investigated that combining binary cross-entropy loss and auxiliary ranking loss improves classification performance. A more recent work Lin et al. (2024) has also shown that the combination of binary cross-entropy loss and ranking loss could address the gradient vanishing issue for negative samples, especially in the context of sparse positive feedback scenarios, therefore improves the classification performance.
@@
While existing literature has shown that the combination of pairwise learning with binary cross-entropy loss in classification problem enhances classification accuracy, only a few He et al. (2024) have studied combining pairwise learning with the regression problem to our knowledge, yet it mainly focus in uplift modeling instead regression problem. Meanwhile, some recent studiesDai et al. (2023); Zhou et al. (2024); Zha et al. (2023); Cheng et al. (2024) try to employ contrastive learning to enhance regression tasks by improving feature representations through leveraging sample similarities and preserving label distance relationships within the feature space. However, these researches are mainly concentrated in the field of image processing instead of proposing generalized methods for tabular data and other real-world tasks.
@@
Furthermore, noise and uncertainty in the dataset adversely affect the regression algorithm’s learning process, leading to decrease of accuracy and robustness, yet less attention is paid to the issue. Uncertainty estimation is the process of quantifying the uncertainty associated with predictions made by machine learning models. Traditional pointwise predictions provide a single value as an estimate without conveying the range of possible outcomes or the likelihood of those outcomes, whereas a probabilistic predictions predict a distribution of possible values, enabling better decisions under uncertainty. Early work on uncertainty estimation formalizes the task as a Bayesian problem by assigning a prior distribution to the parameters of the statistical model and updates the distribution of the parameters conditional on the data to estimate predictive uncertaintyBernardo & Smith (2009). Some approximated Bayesian methodsMackay (1992); Neal (2012); Blundell et al. (2015); Graves (2011) are proposed to alleviate the computation bottleneck. Machine learning models such as generalized additive modelsRigby & Stasinopoulos (2005), random forestsMeinshausen & Ridgeway (2006) and boosting algorithmsHofner et al. (2014); Duan et al. (2020) can also predict specific quantiles of the predictive distribution accurately using quantile regressionMeinshausen & Ridgeway (2006); Bhat et al. (2015) or distribution regressionSchlosser et al. (2019); Thomas et al. (2018). The ensemble methodsHall & Mitchell (2007); Kapetanios et al. (2015); Bassetti et al. (2018) have been shown to be effective in uncertainty estimation or density estimation. Monte Carlo dropoutGal & Ghahramani (2016) is also proposed to estimate predictive uncertainty. The presence of aleatoric uncertainty, as we will demonstrate later, would also affect the learning process of regular pairwise learning.
@@
In this paper, we address the challenge of improving regression tasks by integrating pairwise learning with uncertainty learning using deep probabilistic models. Specifically, we formulate the regression problem as a combination of original regression task together with an auxiliary uncertainty-based learning task; not only does the model generate a more accurate point-wise prediction, but also predicts the relative differences between pairs of data points with the uncertainty associated with these predictions. This formulation allows the model to capture the underlying relationships and dependencies within the data more effectively and provide more reliable and robust predictions in a wide range of regression tasks.
@@
The motivation for our research stems from the limitations of traditional regression methods, particularly in the handling of complex and high-dimensional data and the quantification of uncertainty. We propose a new method called Adaptive Pairwise Regression Learning (AdaPRL). By integrating uncertainty learning, our method can estimate the aleatoric uncertainty level for each sample, allowing the our pairwise regression learning framework to focus more on sample pairs with higher certainty instead of unreliable data points. We also propose adaption of AdaPRL so that it could work not just with single output prediction but also multi-task learning and multivariate time series prediction. Our research has conducted extensive comparative experiments on different regression datasets in wide range of fields including but not limited to recommendation system, time-series forecasting, computer vision, financial and industrial area, which shows that our proposed AdaPRL framework not only improves the accuracy and robustness of the original model but also enhances the ranking ability of the model.
@@
    We conduct extensive experiments on a variety of regression datasets, including image datasets (e.g., AgeDB and CLAP2016), time series datasets (e.g., ETT and Electricity), natrue language understanding datasets (e.g., Jigsaw), recommendation system datasets (e.g., MovieLens-1M and KuaiRec), financial dataet (e.g., Janestreet) and a large-scale proprietary industrial data from a major music streaming app. Extensive experiment results demonstrate that our proposed method can be combined with various backbone models and outperforms the original regression loss on different datasets. Further experiment in the Appendix shows that our method also has the potential to increase generalization ability, robustness and interpretability of modern regression model. Our proposed AdaPRL can also be well integrated with existing method such as RankSimGong et al. (2022) and Ordinal EntropyZhang et al. (2023).
@@ 2.2 Adaptive Pairwise Regression Learning (AdaPRL)
In this session, we analyze the regression problem from a pairwise learning perspective and propose an innovative pairwise regression learning framework called adaptive pairwise regression learning (AdaPRL), where the regression models not only predict the target of the data points but also learn the relative rankings of the sample pairs. This helps the regression model to be aware of the prediction relativity, thus improving the accuracy as well as ranking ability. In order to conduct a thorough analysis of the problem, we address three key issues: the definition of positive and negative pairs, the optimization process, the impact of data noise on pairwise learning, and strategies to mitigate the noise effect.
@@ 2.2.1 Definition of the positive and negative examples
Unlike classification problems, the way to define positive and negative sample pairs has been fully studied in previous workLi et al. (2015); Yan et al. (2022); Yue et al. (2022); Sheng et al. (2023). For example, in the binary classification scenario, the positive or negative nature of samples is typically defined directly by their binary labels. We want the model prediction for positive samples to be higher than for negative samples. Therefore, when one sample is of the positive class and the other is of the negative class, we can naturally expect the model’s predicted score for the positive class sample is higher than that for the negative class sampleLi et al. (2015). In multiclass settings, a prevalent approach for defining positive and negative samples in contrastive learning involves designating samples from the same class or identity as positive pairs and samples from different classes or identities as negative pairsRadford et al. (2021), which facilitates the model’s ability to learn discriminative representations by minimizing the distance between positive pairs while maximizing the distance between negative pairs. Meanwhile, data augmentation and transformation are also used to define positive and negative samples in previous work of contrastive learningChen et al. (2020); He et al. (2020); Grill et al. (2020); Zha et al. (2023).
However, in the settings of regression, the definition of positive and negative samples is not as straightforward. For an unbounded, continuous set of real-number labels for regression, it is difficult to directly define positive and negative samples based on the absolute magnitude of the labels for the reason that a sample with a large magnitude might still be relatively small when compared to another random sample, therefore, we cannot define positive and negative samples based on the absolute values of unbounded continuous labels. Instead, we took an alternative approach to define positive-negative sample pairs based on the relative magnitude of the sample labels. Specifically, for a given pair of samples $(i,j)$, we first calculate difference of the labels :
@@
We follow Li et al. (2015); Tagami et al. (2013) to incorporate hinge loss with pairwise learning. For convenience, assume a threshold parameter $\theta$ and define a pair of samples as a positive-negative pair if and only if the difference between the labels is greater than or equal to the threshold $\theta$ and define a boolean matrix $M$ with size $i\times j$ to represent if $\Delta y_{ij}$ is greater than $\theta$:
@@ 2.2.2 Optimization with auxiliary pairwise ranking loss in regression task
To optimize the model based on the auxiliary pairwise ranking loss (PRL), we need to define a pairwise ranking loss function. Similar to a recent work on uplift modeling with pairwise learningHe et al. (2024), a straightforward method is to make the difference between two samples’ predictions as close as possible to the difference of their labels. To achieve that, first calculate the difference of model prediction of an arbitrary data pair of sample $i$ and sample $j$ as:
@@
Secondly, calculate pairwise loss $L_{prl}$ according to our predefined loss type. Since batch gradient descent is commonly used in practice, we focus on pairwise ranking learning with data batches. Similar to standard regression tasks, where mean squared error (MSE) and mean absolute error (MAE) have been widely applied, the optimization in pairwise regression framework can also be formalized by treating all errors either linearly or quadratically. For simplicity, we refer to the former as the MAE type and the latter as the RMSE type.
@@
where $B$ is mini-batch size, and $D$ is the number of true elements of the hinge matrix M.
@@ 2.2.3 AdaPRL: Tackling noise in pairwise learning by uncertainty estimation
Aleatoric uncertainty is inherent and represents a common challenge in real-world applications, which can arise from ambiguities in the data or labels, errors in data annotation and collection, intrinsic randomness of the data, or random errors from sensors or cameras used for data collection. While the pairwise learning loss proposed above can contribute to improving the accuracy the ranking ability of regression models, which we will demonstrated later in the experimental section, it is important to acknowledge that this straightforward approach has a major drawback when data noise is taken into account. From a distributional perspective of regression tasks, the target $y$ can be considered as a sample from the dataset $D$, where $y$ follows the conditional probability distribution $p(y|x;D)$. Consequently, for any pair of samples $(i,j)$, if any of the samples has a high probability of containing data noise, the ranking and relative magnitude of this pair are also likely to be noisy. Forcing the difference in their predicted values to equal the label difference $\Delta y_{ij}$ during training, as is described above, could negatively impact the precision of both the ranking and regression tasks, exacerbate model overfitting, and can even cause mode collapse.
To further illustrate the potential causes of this phenomenon, consider the following example in which noise cannot be ignored. If both samples $i$ and $j$ have errors in their data annotation or collection, which we annotated as $\epsilon_{i}$ and $\epsilon_{j}$, resulting to sample $i$’s ground truth label $y_{i}$ being altered to $y^{*}_{i}=y_{i}+\epsilon_{i}$ and sample $j$’s label $y_{j}$ being altered to $y^{*}_{j}=y_{j}+\epsilon_{j}$, we then obtain a noise-considering formula of delta label of sample pairs:
@@
where $\sum\epsilon_{ij}$ represents the sum of $\epsilon_{i}$ and $\epsilon_{j}$. Note that the ground truth label of the data is theoretically impossible to obtain because of the uncertain noise, therefore it is evident that the noise of each two samples will cause error accumulation, making it more difficult to predict. As consequence, if we continue to use original form of $L_{prl}$, take MAE type as example, we then obtain:
@@
Assume the noise of data point’s label in $D$ follows a normal distribution $Z\sim\mathcal{N}(\mu,\sigma^{2})$, it is evident that the accumulated error $(\epsilon_{i}\ +\epsilon_{j})$ can be viewed as a random variable follow a different normal distribution.
Therefore,the training process for minimizing $L_{prl-{noise}}$ via gradient descent could be highly unstable because of the randomness induced by distribution sampling. Moreover, in some extreme cases where the noise becomes a dominant term, $\Delta y_{ij}$ and $\Delta y^{*}_{ij}$ could possibly have opposite signs. In that case, the original pairwise regression loss will force the model to predict $\Delta s_{ij}$ as close as possible to $\Delta y^{*}_{ij}$ rather than $\Delta y_{ij}$, causing the model to optimize towards a totally opposite direction.
@@
We integrate a deep probabilistic model into the AdaPRL framework. As discussed above, we assume that the label follows a Gaussian distribution $Z\sim\mathcal{N}(\mu,\sigma^{2})$. Firstly, we define an auxiliary deep probabilistic model with NLL loss and let the model predict the mean $\mu_{i}$ together with the associated uncertainty $\sigma_{i}$, of each sample $s_{i}$.
Secondly, for every pair of sample $(i,j)$, we measure the variation of the distribution of $(y_{i}-y_{j})$. The distribution of $(y_{i}-y_{j})$ will follow another Gaussian distribution:
@@
Since both sample are derived independently from the dataset, we assume that two Gaussian distribution are independent for simplification. As a result, the distribution of $(y_{i}-y_{j})$ can be simplified to:
@@
because the covariance between $y_{i}$ and $y_{j}$ equals to 0.
Thirdly, we can generate a uncertainty matrix $U$ with size $i\times j$ to represent the aleatoric uncertainty of data pairs $(i,j)$:
@@
In the fourth step, conduct a inverse min-max transformation of matrix $U$ to generate a normalized matrix, which is called the confidence matrix $C$:
@@
where $max(U)$ and $min(U)$ indicate the maximum and minimum value in matrix $U$ respectively. An additional weight factor of 2 is included to empirically rescale the matrix. After transformation, the value of $C_{ij}$ is greater if the uncertainty $U_{ij}$ is smaller and vice versa, which means that the normalized matrix $C$ indicates the confidence of data pairs. The larger $C_{ij}$, the more confidence we have in the joint probability distribution of data pairs $(i,j)$ and their relative difference.
Finally, we combine the confidence matrix $C$ with the original pairwise regression learning loss, then obtain pairwise regression loss with confidence (CPRL) :
@@
where $L_{reg}$ could be any arbitrary regression loss function such as mean squared error (MSE), mean absolute error (MAE), Huber loss. $\alpha$ is the hyperparameter to control the weight of pairwise regression loss with confidence. Complete process of the AdaPRL algorithm is described in 1.

### v3

@@ Abstract
Current deep regression models usually learn in a point-wise way that treats each sample as an independent input, neglecting the relative ordering among different data. Consequently, the regression model could neglect the data’s interrelationships, potentially resulting in suboptimal performance. Moreover, the existence of aleatoric uncertainty in the training data may drive the model to capture non-generalizable patterns, contributing to increased overfitting. To address these issues, we propose a novel adaptive pairwise learning framework for regression tasks (AdaPRL) which leverages the relative differences between data points and integrates with deep probabilistic models to quantify the uncertainty associated with the predictions. Additionally, we adapt AdaPRL for applications in multi-task learning and multivariate time series forecasting. Extensive experiments with several real-world regression datasets including recommendation systems, age prediction, time series forecasting, natural language understanding, finance, and industry datasets show that AdaPRL is compatible with different backbone networks in various tasks and achieves state-of-the-art performance on the vast majority of tasks without extra inference cost, highlighting its notable potential including enhancing prediction accuracy and ranking ability, increasing generalization capability, improving robustness to noisy data, improving resilience to reduced data, and enhancing interpretability. Experiments also show that AdaPRL can be seamlessly incorporated into recently proposed regression frameworks to gain performance improvement.
@@ 1 Introduction
In recent years, regression tasks have become increasingly important in various domains, including computer vision, time series forecasting, recommendation systems, finance, and so on. Traditional regression methods which integrate with point-wise loss functions such as mean squared error (MSE), mean absolute error (MAE) and Huber have been widely adopted due to their simplicity and effectiveness. However, these traditional pointwise methods often struggle with capturing the nuanced relationships between data points, potentially resulting in less effective performance. Pairwise learning, on the other hand, offers a promising alternative by focusing on relative differences between pairs of data points rather than on absolute values.
@@
To overcome the limitations of point-wise learning, Learning-to-Rank (LTR) approaches have been introduced, where pairwise methods aim to ensure that the predicted score of a positive example surpasses that of a negative sample for each positive-negative pair, and listwise methods promote higher rankings for positive samples across the entire list of items. Representative pairwise learning methods include RankSVM Joachims (2002), GBRank Zheng et al. (2007), RankNet Burges et al. (2005), LambdaRank Burges et al. (2006) and LambdaMART Burges (2010). Some previous studies Li et al. (2015); Yan et al. (2022); Yue et al. (2022); Sheng et al. (2023) have investigated that combining binary cross-entropy loss and auxiliary ranking loss improves classification performance. A more recent work Lin et al. (2024) has also shown that the combination of binary cross-entropy loss and ranking loss can address the gradient vanishing issue for negative samples, especially in the context of sparse positive feedback scenarios, thereby improving the classification performance.
@@
While existing literature has shown that the combination of pairwise learning with binary cross-entropy loss in classification problems enhances classification accuracy, only a few studies He et al. (2024) have studied combining pairwise learning with the regression problem to our knowledge, yet they mainly focus on uplift modeling rather than on regression problems. Meanwhile, some recent studies Dai et al. (2023); Zhou et al. (2024); Zha et al. (2023); Cheng et al. (2024) try to employ contrastive learning to enhance regression tasks by improving feature representations through leveraging sample similarities and preserving label distance relationships within the feature space. However, these researches are mainly concentrated in the field of image processing instead of proposing generalized methods for tabular data and other real-world tasks.
@@
Furthermore, noise and uncertainty in the dataset adversely affect the regression algorithm’s learning process, leading to a decrease of accuracy and robustness, yet less attention is paid to the issue. Uncertainty estimation is the process of quantifying the uncertainty associated with the predictions made by machine learning models. Traditional pointwise predictions provide a single value as an estimate without conveying the range of possible outcomes or their likelihood, whereas probabilistic predictions predict a distribution of possible values, enabling better decisions under uncertainty. Early work on uncertainty estimation formalizes the task as a Bayesian problem by assigning a prior distribution to the parameters of the statistical model and updating the distribution of the parameters conditional on the data to estimate predictive uncertainty Bernardo & Smith (2009). Some approximated Bayesian methods Mackay (1992); Neal (2012); Blundell et al. (2015); Graves (2011) are proposed to alleviate the computation bottleneck. Machine learning models such as generalized additive models Rigby & Stasinopoulos (2005), random forests Meinshausen & Ridgeway (2006) and boosting algorithms Hofner et al. (2014); Duan et al. (2020) can also predict specific quantiles of the predictive distribution accurately using quantile regression Meinshausen & Ridgeway (2006); Bhat et al. (2015) or distribution regression Schlosser et al. (2019); Thomas et al. (2018). The ensemble methods Hall & Mitchell (2007); Kapetanios et al. (2015); Bassetti et al. (2018) have been shown to be effective in uncertainty estimation or density estimation. Monte Carlo dropout Gal & Ghahramani (2016) is also proposed to estimate predictive uncertainty. The presence of aleatoric uncertainty, as we will demonstrate later, also affects the learning process of regular pairwise learning.
@@
In this paper, we address the challenge of improving regression tasks by integrating pairwise learning and uncertainty learning using deep probabilistic models. Specifically, we formulate the regression problem as a combination of the original regression task together with an auxiliary uncertainty-based learning task; not only does the model generate a more accurate point-wise prediction, but it also predicts the relative differences between pairs of data points with the uncertainty associated with these predictions. This formulation allows the model to capture the underlying relationships and dependencies within the data more effectively and provides more reliable and robust predictions in a wide range of regression tasks.
@@
The motivation for our research stems from the limitations of existing regression methods, especially in handling complex and high-dimensional data, as well as quantifying the uncertainty of the data. We propose a new method called Adaptive Pairwise Regression Learning (AdaPRL). By integrating uncertainty learning, our method can estimate the aleatoric uncertainty level for each sample, allowing our pairwise regression learning framework to focus more on sample pairs with higher certainty instead of unreliable data points. We also propose adaptation of AdaPRL so that it could work not just with single output prediction but also multi-task learning and multivariate time series prediction. Our research has conducted extensive comparative experiments on different regression datasets in a wide range of fields including but not limited to recommendation systems, time-series forecasting, computer vision, and financial and industrial areas, which shows that our proposed AdaPRL framework not only improves the accuracy and robustness of the original model but also enhances the ranking ability of the model.
@@
    We conduct extensive experiments on a variety of regression datasets, including image datasets (e.g., AgeDB and CLAP2016), time series datasets (e.g., ETT and Electricity), natrue language understanding datasets (e.g., Jigsaw), recommendation systems datasets (e.g., MovieLens-1M and KuaiRec), financial dataet (e.g., Janestreet) and a large-scale proprietary industrial data from a major music streaming app. Extensive experiment results demonstrate that our proposed method can be combined with various backbone models and outperforms the original regression loss on different datasets. Further experiment in the Appendix shows that our method also has the potential to increase generalization ability, robustness and interpretability of modern regression model. Our proposed AdaPRL can also be well integrated with existing regression methods such as RankSim Gong et al. (2022) and Ordinal Entropy Zhang et al. (2023).
@@ 2.2 Adaptive Pairwise Regression Learning (AdaPRL)
In this session, we analyze the regression problem from a pairwise learning perspective and propose an innovative pairwise ranking framework for regression called Adaptive Pairwise Regression Learning (AdaPRL), where the regression models not only predict the target of the data points but also learn the relative rankings and certainty of the sample pairs. This helps the regression model to be aware of the prediction relativity, thus improving the accuracy as well as the ranking ability. To conduct a comprehensive analysis of the problem, we address several key issues: the definition of positive and negative pairs, the optimization process, the impact of data noise on pairwise learning, and how our proposed method mitigates this noise effect.
@@ 2.2.1 Definition of the positive and negative examples
Unlike classification problems, the way to define positive and negative sample pairs has been fully studied in previous work Li et al. (2015); Yan et al. (2022); Yue et al. (2022); Sheng et al. (2023). For example, in the binary classification scenario, the positive or negative nature of samples is typically defined directly by their binary labels. We want the model prediction for positive samples to be higher than for negative samples. Therefore, when one sample is of the positive class and the other is of the negative class, we can naturally expect the model’s predicted score for the positive class sample to be higher than that for the negative class sample Li et al. (2015). In multi-class classification settings, a prevalent approach for defining positive and negative samples in contrastive learning involves designating samples from the same class or identity as positive pairs and samples from different classes or identities as negative pairs Radford et al. (2021), which facilitates the model’s ability to learn discriminative representations by minimizing the distance between positive pairs while maximizing the distance between negative pairs. Meanwhile, data augmentation and transformation techniques are frequently used to define positive and negative samples in earlier research on contrastive learning Chen et al. (2020); He et al. (2020); Grill et al. (2020); Zha et al. (2023).
However, the definition of positive and negative samples is not as straightforward in the settings for regression tasks. In the context of regression problems, which typically involve predicting unbounded and continuous labels, it is difficult to directly define positive and negative samples solely based on the absolute magnitude of the labels because a sample with a large magnitude might still be relatively small when compared to another sample; therefore, we took an alternative approach to define positive-negative sample pairs based on the relative magnitude of the samples’ labels.
Specifically, for a given pair of samples $(i,j)$, first we need to calculate the difference of labels:
@@
We follow Li et al. (2015); Tagami et al. (2013) to incorporate hinge loss with pairwise learning. For convenience, assume a threshold parameter $\theta$ and define a pair of samples as a positive-negative pair if and only if the difference between the labels is greater than or equal to the threshold $\theta$ and define a boolean matrix $M$ with size $i\times j$ to represent if $\Delta y_{ij}$ is greater than $\theta$, that is to say:
@@ 2.2.2 Optimization with auxiliary pairwise ranking loss in regression task
To optimize the model based on the auxiliary pairwise ranking loss (PRL), we need to define a pairwise ranking loss function. Similar to a recent work on uplift modeling with pairwise learning He et al. (2024), a straightforward way of combining pairwise learning with regression tasks is to make the difference between two samples’ predictions as close as possible to the difference of their labels. To achieve that, firstly calculate the difference of the model prediction of an arbitrary data pair of sample $i$ and sample $j$ as:
@@
Secondly, calculate pairwise loss $L_{prl}$ according to the predefined loss type. Since batch gradient descent is commonly used in practice, we focus on pairwise ranking learning with data batches. Similar to standard regression tasks, where mean squared error (MSE) and mean absolute error (MAE) have been widely applied, the optimization target for the pairwise regression framework can also be formalized by treating all errors either linearly or quadratically. We refer to the former as the MAE type and the latter as the RMSE type for the purpose of simplicity.
Specifically, if MAE type pairwise learning loss is used, then pairwise regression loss becomes:
@@
where $B$ is mini-batch size, and $D$ is the number of true elements of the hinge matrix M. We will refer to the method as pairwise regression learning for brevity in subsequent sections.
@@ 2.2.3 AdaPRL: Tackling noise in pairwise learning by uncertainty estimation
Aleatoric uncertainty is inherent and represents a common challenge in real-world applications, which can arise from ambiguities in the data or labels, errors in data annotation and collection, intrinsic randomness of the data, or random errors from sensors or cameras used for data collection Hüllermeier & Waegeman (2021). While the standard pairwise learning loss proposed above contributes to improving the accuracy and the ranking ability of regression models, which we will demonstrate in the experiments, it is important to acknowledge that this straightforward approach has a major drawback when data noise or uncertainty is taken into account.
From a distributional perspective of regression tasks, the target $y$ can be considered as a random sample drawn from the dataset $D$, where $y$ follows the conditional probability distribution $p(y|x;D)$. Consequently, for any pair of samples $(i,j)$ that has a high probability of containing data noise, the ranking and relative magnitude of this pair are also likely to be noisy. Therefore, forcing the difference of their predicted values to be equal to the label difference $\Delta y_{ij}$ during training, as described above, could negatively impact the precision of both the ranking and regression tasks, exacerbate model overfitting, or even cause mode collapse.
To further explain the phenomenon, consider the following example where random noise cannot be ignored. Assume that both samples $i$ and $j$ contain a certain level of noise during the data annotation or collection phase, which we annotated as $\epsilon_{i}$ and $\epsilon_{j}$, resulting in sample $i$’s ground truth label $y_{i}$ being altered to $y^{*}_{i}=y_{i}+\epsilon_{i}$ and sample $j$’s label $y_{j}$ being altered to $y^{*}_{j}=y_{j}+\epsilon_{j}$. We then obtain a noise-considering formula of delta label of sample pairs:
@@
where $\sum\epsilon_{ij}$ represents the sum of $\epsilon_{i}$ and $\epsilon_{j}$. Note that the ground truth label of the data is theoretically impossible to obtain because of the aleatoric uncertainty; therefore, it is evident that operation 6 will cause error accumulation, making it more difficult to learn. As a consequence, if we continue to use the original form of $L_{prl}$, take MAE type as an example without loss of generality, we obtain:
@@
Assume the noise of the label in $D$ follows a Gaussian distribution $Z\sim\mathcal{N}(\mu,\sigma^{2})$, it is obvious that the accumulated error $(\epsilon_{i}\ +\epsilon_{j})$ can also be viewed as a random variable following another Gaussian distribution with higher variance.
Therefore, the training process of minimizing $L_{prl-{noise}}$ via gradient descent could be highly unstable because of the randomness induced by distribution sampling. Moreover, in some extreme cases where the noise becomes a dominant term, $\Delta y_{ij}$ and $\Delta y^{*}_{ij}$ could possibly have opposite signs. In that case, the original pairwise regression loss will force the model to predict $\Delta s_{ij}$ as close as possible to $\Delta y^{*}_{ij}$ rather than $\Delta y_{ij}$, causing the model to optimize towards a totally opposite direction.
@@
We integrate a deep probabilistic model into the AdaPRL framework. As discussed above, assume that the label follows a Gaussian distribution $Z\sim\mathcal{N}(\mu,\sigma^{2})$. Firstly, we define an auxiliary deep probabilistic model with NLL loss and make the model predict the mean $\mu_{i}$ together with the associated uncertainty $\sigma_{i}$ for each sample $s_{i}$.
Secondly, for every pair of sample $(i,j)$, we measure the variation of the distribution of $(y_{i}-y_{j})$. The distribution of $(y_{i}-y_{j})$ will follow the following Gaussian distribution:
@@
Since both samples are derived independently from the dataset, for simplification, we assume that two Gaussian distributions are independent. As a result, the distribution of $(y_{i}-y_{j})$ can be simplified to:
@@
because the covariance between $y_{i}$ and $y_{j}$ equals 0.
Thirdly, we can generate an uncertainty matrix $U$ with size $i\times j$ to represent the aleatoric uncertainty of pair $(i,j)$:
@@
In the fourth step, conduct an inverse min-max transformation of matrix $U$ to generate a normalized matrix, which is called the confidence matrix $C$:
@@
where $max(U)$ and $min(U)$ indicate the maximum and minimum value in matrix $U$ respectively. An additional weight factor of 2 is included to empirically rescale the matrix. After the transformation, the value of $C_{ij}$ is greater if the uncertainty $U_{ij}$ is smaller and vice versa, which means that the normalized matrix $C$ indicates the confidence of data pairs. The larger $C_{ij}$, the more confidence we have in the joint probability distribution of data pairs $(i,j)$ and their relative difference.
Finally, combine the confidence matrix $C$ with the original pairwise regression learning loss, then we would obtain pairwise regression loss with confidence (CPRL) :
@@
where $L_{reg}$ could be any arbitrary regression loss function such as mean squared error (MSE), mean absolute error (MAE), or Huber loss. Hyperparameter $\alpha$ is a float to control the weight of pairwise regression loss with confidence. The complete AdaPRL algorithm is described in 1.